1. The table shows the jobs to be scheduled and estimated process times for the jobs. The promised delivery times are shown in the table.

Job process time delivery due time slack time
A 4 75
B 16 90
C 2 30
D 27 90
E 5 90
F 38 75
G 26 120

Determine the average process completion time and average delay using (i) the shortest processing time first and (ii) the minimum slack time first dispatch rules.
Answer #1

Here, Important Formulas are

slack time Remaining = Delivery Due time - Process Time

Flow Time = cumulative process time,

Lateness = Flow time - Delivery Due time,

Average process completion time = Total Flow time/ Number of Jobs

Average Delay or lateness = Total Lateness/number of Jobs

For Shortest processing time, we will consider the activities in order of shortest processing time,

For Minimum Slack time, we we will consider the activities in order of Minimum slack time,

1 B C D E F
2 Job process time delivery due time slack time Remaining = Delivery Due time - Process Time
3 A 4 75 71
4 B 16 90 74
5 C 2 30 28
6 D 27 90 63
7 E 5 90 85
8 F 38 75 37
9 G 26 120 94
10
11 Shortest Processing Time
12 Job Process time Flow Time delivery due time Lateness
13 C 2 2 30 0
14 A 4 6 75 0
15 E 5 11 90 0
16 B 16 27 90 0
17 G 26 53 120 0
18 D 27 80 90 0
19 F 38 118 75 43
20 Average Process completion Time Average Lateness
21 42.43 6.14
22
23 Minimum Slack Time
24 Job Process time Flow Time delivery due time Lateness
25 C 2 2 30 0
26 F 38 40 75 0
27 D 27 67 90 0
28 A 4 71 75 0
29 B 16 87 90 0
30 E 5 92 90 2
31 G 26 118 120 0
32 Average Process completion Time Average Lateness
33 68.14 0.29

Formula

1 B C D E F
2 Job process time delivery due time slack time Remaining = Delivery Due time - Process Time
3 A 4 75 =D3-C3
4 B 16 90 =D4-C4
5 C 2 30 =D5-C5
6 D 27 90 =D6-C6
7 E 5 90 =D7-C7
8 F 38 75 =D8-C8
9 G 26 120 =D9-C9
10
11 Shortest Processing Time
12 Job Process time Flow Time delivery due time Lateness
13 C 2 =C13 30 0
14 A 4 =C14+D13 75 0
15 E 5 =C15+D14 90 0
16 B 16 =C16+D15 90 0
17 G 26 =C17+D16 120 0
18 D 27 =C18+D17 90 0
19 F 38 =C19+D18 75 =D19-E19
20 Average Process completion Time Average Lateness
21 =AVERAGE(D13:D19) =AVERAGE(F13:F19)
22
23 Minimum Slack Time
24 Job Process time Flow Time delivery due time Lateness
25 C 2 =C25 30 0
26 F 38 =C26+D25 75 0
27 D 27 =C27+D26 90 0
28 A 4 =C28+D27 75 0
29 B 16 =C29+D28 90 0
30 E 5 =C30+D29 90 =D30-E30
31 G 26 =C31+D30 120 0
32 Average Process completion Time Average Lateness
33 =AVERAGE(D25:D31) =AVERAGE(F25:F31)